### v1

# Beyond the Doors of Perception: Vision Transformers Represent Relations Between Objects

Abstract
Though vision transformers (ViTs) have achieved state-of-the-art performance in a variety of settings, they exhibit surprising failures when performing tasks involving visual relations. This begs the question: how do ViTs attempt to perform tasks that require computing visual relations between objects? Prior efforts to interpret ViTs tend to focus on characterizing relevant low-level visual features. In contrast, we adopt methods from mechanistic interpretability to study the higher-level visual algorithms that ViTs use to perform abstract visual reasoning. We present a case study of a fundamental, yet surprisingly difficult, relational reasoning task: judging whether two visual entities are the same or different. We find that pretrained ViTs fine-tuned on this task often exhibit two qualitatively different stages of processing despite having no obvious inductive biases to do so: 1) a perceptual stage wherein local object features are extracted and stored in a disentangled representation, and 2) a relational stage wherein object representations are compared. In the second stage, we find evidence that ViTs can learn to represent somewhat abstract visual relations, a capability that has long been considered out of reach for artificial neural networks. Finally, we demonstrate that failure points at either stage can prevent a model from learning a generalizable solution to our fairly simple tasks. By understanding ViTs in terms of discrete processing stages, one can more precisely diagnose and rectify shortcomings of existing and future models.
1 Introduction
Despite the well-established successes of transformer models (Vaswani et al., 2017) for a variety of vision applications (ViTs; Dosovitskiy et al. (2020)) – notably image generation and classification – there has been comparatively little breakthrough progress on complex tasks involving relations between visual entities, such as visual question answering (Schwenk et al., 2022) and image-text matching (Thrush et al., 2022; Liu et al., 2023; Yuksekgonul et al., 2022). One fundamental difference between these tasks is that the former is largely semantic – relying on pixel-level image features that correlate with learned class labels – whereas the latter often involves syntactic operations – those which are independent of pixel-level features (Ricci et al., 2021; Hochmann et al., 2021). Though the ability to compute over abstract visual relations is thought to be fundamental to human visual intelligence (Ullman, 1987; Hespos et al., 2021), the ability of neural networks to perform such syntactic operations has been the subject of intense debate (Fodor & Pylyshyn, 1988; Marcus, 2003; Chalmers, 1992; Quilty-Dunn et al., 2023; Lake et al., 2017; Davidson et al., 2024).
Much prior work has attempted to empirically resolve whether or not vision networks can implement an abstract, relational operation, typically by behaviorally assessing the model’s ability to generalize to held-out stimuli (Fleuret et al., 2011; Zerroug et al., 2022; Kim et al., 2018; Puebla & Bowers, 2022; Tartaglini et al., 2023). However, strikingly different algorithms might produce the same model behavior, rendering it difficult to characterize whether models do or do not possess an abstract operation (Lepori et al., 2023a). This problem is exacerbated when analyzing pretrained models, whose opaque training data renders it difficult to distinguish true generalization from memorization (McCoy et al., 2023). In this work, we employ newly-developed techniques from mechanistic interpretability to characterize the algorithms learned by ViTs. Analyzing the internal mechanisms of models enables us to more precisely understand how they attempt to implement relational operations, allowing us to more clearly diagnose problems in current and future models when applied to complex visual tasks.
Figure 1: Two same-different tasks. (a) Discrimination: “same” images contain two objects with the same color and shape. Objects in “different” images differ in at least one of those properties—in this case, both color and shape. (b) RMTS: “same” images contain a pair of objects that exhibit the same relation as a display pair of objects in the top left corner. In the image on the left, both pairs demonstrate a “different” relation, so the classification is “same” (relation). “Different” images contain pairs exhibiting different relations.
One of the most fundamental of these abstract operations is identifying whether two objects are the same or different. This operation undergirds human visual and analogical reasoning (Forbus & Lovett, 2021; Cook & Wasserman, 2007), and is crucial for answering a wide variety of common visual questions, such as “How many plates are on the table?" (as each plate must be identified as an instance of the same object) or “Are Mary and Bob reading the same book?” (Ricci et al., 2021). Indeed, same-different judgments can be found across the animal kingdom, being successfully captured by bees (Giurfa et al., 2001), ducklings (Martinho III & Kacelnik, 2016), primates (Thompson & Oden, 2000), and crows (Cook & Wasserman, 2007).
We analyze ViTs trained on two same-different tasks: an identity discrimination task, which constitutes the most basic instances of the abstract concept of same-different (and is most commonly studied in artificial systems) (Newport, 2021), and a relational match-to-sample task, which requires explicitly representing and manipulating an abstract concept of “same” or “different” (Cook & Wasserman, 2007; Geiger et al., 2023). See Figure 1 for examples of each. We relate the algorithms that models adopt to their downstream behavior, including compositional and OOD generalization [^1] Code is available at an anonymous github repo..
Our main contributions include:
  1. 1.
    Inspired by the infant and animal abstract concept learning literature, we introduce a synthetic visual relation match-to-sample (RMTS) task, which assesses a model’s ability to represent and compute over the abstract concept of “same” or “different”.
  2. 2.
    We identify a processing pipeline within the layers of several – but not all – pretrained ViTs, consisting of a “perceptual” stage followed by a more abstract “relational” stage. We characterize each stage individually, demonstrating that the perceptual stage produces disentangled object representations (Higgins et al., 2018), while the relational stage implements fairly abstract (i.e. invariant to perceptual properties of input images) relational computations – an ability that has been intensely debated since the advent of neural networks (Fodor & Pylyshyn, 1988).
  3. 3.
    We demonstrate that deficiencies in the perceptual stage can completely prevent models from learning abstract relational operations. By rectifying this first stage, these models can solve simple relational operations (i.e., discrimination) yet still struggle to learn more complex operations (i.e., RMTS).
2 Methods
Discrimination Task: We analyze a task involving the most basic instance of the same-different relation: the discrimination task. This task is well studied in machine learning (Kim et al., 2018; Tartaglini et al., 2023; Puebla & Bowers, 2022), and requires a mere comparison between two objects presented in a single image. It is also comparable to Identity-Match-to-Sample and Discrimination procedures in the cognitive sciences (Hochmann et al., 2017). Stimuli in the discrimination task consist of images containing two simple objects (See Figure 1, left). Each object may take on one of 16 different shapes and one of 16 different colors. The model is trained to distinguish whether these two tokens are the same along both color and shape dimensions or different along at least one of these dimensions. We train on 6,400 of these images. Notably, our stimuli are patch-aligned. Vision transformers tokenize images into square patches of N×$\times$×N pixels (in our case, $N=\{16,32\}$). We generate datasets such that objects reside completely within the bounds of a single patch (for the $N=32$ models) or within the bounds of exactly 4 patches (for the $N=16$ models). This simplification allows us to straightforwardly adopt techniques from mechanistic interpretability. To increase the difficulty of the task, tokens are randomly placed within patch boundaries, and Gaussian noise is sampled and added to the tokens.
Relational Match-to-Sample Task: We also analyze a more abstract iteration of the same-different relation using a relational match-to-sample (RMTS) design. In this task, the model must generate an explicit representation of “sameness” and “difference”, and then operate over these representations (Martinho III & Kacelnik, 2016; Hochmann et al., 2017). Notably, though many species can solve the discrimination task, animals struggle to solve the RMTS task (Penn et al., 2008). Indeed, human children fail to solve RMTS tasks until around age 4 and only exhibit robust performance by approximately age 6 (Hochmann et al., 2017; Holyoak & Lu, 2021). Stimuli in the RMTS task contain 4 objects grouped in two pairs (See Figure 1, right). The “display” pair always occupies patches in the image’s top-left. The “sample” pair can occupy any other position. The task is defined as follows: for each pair, produce a same-different judgment (as in the discrimination task). Then, compare these judgments; if both pairs exhibit the same intermediate judgment (i.e., both pairs exhibit “same” or both pairs exhibit “different”), then the label is “same”. Otherwise, the label is different. Objects are identical to those found in the discrimination task, and they are patch-aligned in the same manner.
Figure 2: Attention Pattern Analysis (a) CLIP Discrimination: Top: Heatmap showing the distribution of “local” (blue) vs. “global” (red) attention heads throughout a finetuned CLIP B/16 model. We see that local heads tend to cluster in the first several layers of the model, followed by a transition to global heads. Bottom: Line graph plotting the maximum proportion of attention paid by a single head to patches that are: Within the same object (within-object; WO), Within the other object that is being compared to the current patch’s object patch (within-pair; WP), in the background (BG). We see that WO values peak in early layers, followed by WP values, and finally, BG values. (b) From-Scratch Discrimination: We repeat the analysis in (a). This model does seem to contain local heads. (c) CLIP RMTS: We repeat the analysis for a CLIP model finetuned on RMTS. Top: Our results largely hold from (a). Bottom: We track another attention pattern: attention between pairs of objects (between pair; BP). We find that first, WO peaks, then WP, then BP, and finally BG. This accords with the hierarchical computations implied by the RMTS task. (d) DINO RMTS: We repeat the analysis in (c) for a DINO model and find no such pattern.
Models: Tartaglini et al. (2023) demonstrated that CLIP-pretrained ViTs (Radford et al., 2021) can achieve high performance generalizing the same-different relation to out-of-distribution stimuli when fine-tuned on a same-different discrimination task. Thus, we primarily focus our analysis on this model [^2] Both 16-patch and 32-patch versions. Results for the 32-patch variant are presented in Appendix C.. In later sections, we compare to models pretrained using DINO (Caron et al., 2021) and ImageNet classification (Russakovsky et al., 2015; Dosovitskiy et al., 2020) objectives. We also train a randomly-initialized ViT model on each task. All models are finetuned on either a discrimination or RMTS task for 200 epochs, using the AdamW optimizer with a learning rate of 1E-6. We perform a sweep over learning rate schedulers (Exponential with a decay rate of $0.95$ and ReduceLROnPlateau with a patience of $40$). Models are selected by validation accuracy, and test accuracy is reported. Our results affirm and extend those presented in Tartaglini et al. (2023) – CLIP-pretrained ViTs perform extremely well on both tasks, achieving $\geq$97accuracy on a held-out test set. See Appendix B for results from all models.
3 Two-Stage Processing in ViTs
We now begin to characterize the internal mechanisms underlying the success of the CLIP ViT models. We discover that processing occurs in two distinct stages within the model: a perceptual stage, where the object tokens strongly attend to other tokens within the same object, and a relational stage, where tokens in one object attend to tokens in another object (or pair of objects).
Methods – Attention Pattern Analysis: We explore the operations performed by the model’s attention heads, which “read” from particular patches and “write” that information to other patches (Elhage et al., 2021). In particular, we are interested in the flow of information within individual objects, between the two objects, and (in the case of RMTS) between two pairs of objects. We refer to attention heads that consistently exhibit within-object patterns across images as local attention heads and heads that attend to other tokens global attention heads, respectively. To classify an attention head as local or global, we score the head from 00 to $1$, where values closer to 00 indicate local operations and values closer to $1$ indicate global operations. To compute the score for an individual head, we collect attention patterns from the head on $500$ randomly selected “same” and “different” images ($1,000$ images total). Then, for each object in a given image, we compute the proportion of attention from the object’s patches to patches that do not comprise the object (excluding the CLS token) — including patches containing the other object(s) and non-object background tokens. [^3] We include attention from object to non-object tokens because we observe that models often move object information to a set of background register tokens (Darcet et al., 2023). See Appendix D. This yields two proportions, one for each object in the image. The attention head’s score for the image is the maximum of these two proportions. Finally, these scores are averaged across the images to produce the final score.
Results: Attention head scores for CLIP fine-tuned on the discrimination and RMTS tasks are displayed in the heatmaps of Figure 2(a) and (c), respectively. The first half of these models is dominated by attention heads that most often perform local operations (blue cells). See Appendix D for examples of attention patterns. In the intermediate layers, attention heads begin to perform global operations reliably, and the deeper layers of the model are dominated by global heads. The prevalence of these two types of operations clearly demarcates two processing stages in CLIP ViTs: a perceptual stage where within-object processing occurs, followed by a relational stage where between-object processing occurs. These stages are explored in further depth in Sections 4 and  5 respectively [^4] We note that this attention pattern analysis is conceptually similar to Raghu et al. (2021), which demonstrated a general shift from local to global attention in ViTs. However, our analysis defines local vs. global heads in terms of objects rather than distances between patches..
Figure 3: (a) Interchange interventions The base image exhibits the “different” relation, as the two objects differ in color. An interchange intervention extracts color information from a the intermediate representations generated by the same model run on a different image (the source image), and patches color information from the source image into the model’s intermediate representations of the base image. If successful, the intervened model will now return “same” when run on the base image. DAS is optimized to succeed at interchange interventions. (b) Disentanglement Results: We report the success of interchange interventions on shape and color across layers for CLIP B/16 finetuned on either the discrimination or RMTS task. We find that these properties are disentangled early in the model – one property can be manipulated without interfering with the other. The background is colored according to the heatmap in Figure 2(a).
The line charts in Figure 2(a) and (b) show maximal scores of each attention head type (local and global) in each layer. Since (by construction) values closer to 00 indicate local heads, we plot these values subtracted from $1$. The global attention heads are further broken down into two subcategories for the discrimination task: object attention heads, whereby the tokens of one object attend to tokens associated with the object it is being compared to, and other attention heads whereby object tokens attend to background tokens. We add a third subcategory for RMTS: pair attention heads attend to tokens in the other pair of objects (e.g., a display object attending to a sample object). For both tasks, objects strongly attend to themselves during the first six layers, with a peak at layer 3. Throughout this perceptual stage, object heads steadily increase in prominence and peak in layer 6 (discrimination) or 5 (RMTS). In RMTS models, this object peak is followed by a pair peak, recapitulating the expected sequence of steps that one might use to solve RMTS. Notably, the object (and pair) peaks occur precisely where an abrupt transition from perceptual operations to relational operations occurs. Around layer 4, object attention to a set of background tokens begins to increase; after layer 6, object-to-background attention accounts for nearly all outgoing attention from object tokens. This suggests that processing may have moved into a set of register tokens (Darcet et al., 2023).
Notably, this two-stage processing pipeline is not trivial to learn – neither a randomly initialized model trained on the discrimination task nor a DINO model trained on RMTS (Figure 2(b) and (d)) exhibit any obvious transition from local to global operations. We note that this two-stage processing pipeline loosely recapitulates the processing sequence found in biological vision systems: image representations are first formed during a feedforward sweep of the visual cortex, then feedback connections enable relational reasoning over these representations Kreiman & Serre (2020).
4 The Perceptual Stage
Attention between tokens is largely restricted to other tokens within the same object in the perceptual stage, but to what end? We find that these layers produce disentangled local object representations which encode shape and color. These properties are represented in separate linear subspaces within the intermediate representations of CLIP-pretrained ViTs.
Methods – Distributed Alignment Search: Distributed Alignment Search (DAS) is used to identify whether particular variables are causally implicated in a model’s computation (Geiger et al., 2024; Wu et al., 2024c) [^5] DAS has recently come under scrutiny for being too expressive (and thus not faithful to the model’s internal algorithms) when misused (Makelov et al. (2023), cf. Wu et al. (2024b)) Following best practices, we deploy DAS on the residual stream. We use the pyvene library for all DAS experiments (Wu et al., 2024a).. Given a neural network $M$, hypothesized high-level causal model $C$, and high-level variable v, DAS attempts to isolate a linear subspace $s$ of the residual stream states generated by $M$ that represents $v$ (i.e. $s$ takes on a value $s_{1}$ to represent $v_{1}$, $s_{2}$ to represent $v_{2}$, and so on). The success of DAS is measured by the success of counterfactual interventions. If $C(v_{1})=y_{1}$ and $C(v_{2})=y_{2}$, and $M(x)=y_{1}$ for some input $x$, does replacing $s_{1}$ with $s_{2}$ change the model’s decision to $y_{2}$?
Figure 4: (a) Novel Representations Analysis: Using trained DAS interventions, we can inject any vector into the shape or color subspaces that a model produces, allowing us to test whether the same-different operation can be computed over arbitrary vectors. We intervene on a “different”’ image – which differs only in its color property – by patching a novel color (an interpolation of red and black) into both objects in order to flip the decision to “same”. (b) Results: Results from performing the novel representations analysis using four methods for generating novel representations: adding observed representations, interpolating observed representations, per-dimension sampling using a distribution derived from observed representations, and finally sampling randomly from a mean 0, standard deviation 1 normal distribution. We see that the model’s same-different operation seems to generalize to vectors generated by adding (and to generalize somewhat to interpolated vectors) in early layers but not to sampling. The background is colored according to the heatmap in Figure 2(a).
[…]
Results: We identify independent linear subspaces for color and shape in the intermediate representations produced in the early layers of a CLIP-pretrained ViT (See Figure3(b)). In other words, we can extract either color or shape information from one object and inject into another object. This holds for both discrimination and RMTS tasks. One can conclude that at least one function of the perceptual stage is to form disentangled local object representations, which are then used to solve same-different tasks. Notably, we see that local object representations are formed in the first few layers, becoming increasingly irrelevant deeper into the model. Indeed, intervening on the intermediate representations of object tokens at layers 5 and beyond results in chance intervention performance or worse.
5 The Relational Stage
We now characterize the relational stage, where attention between tokens largely excludes other tokens within the same object. We hypothesize that this stage takes in the object representations formed by the perceptual stage and computes relational same-different operations over them. We find that the operations implemented by these relational layers are somewhat abstract in that 1) they do not rely on memorizing individual objects and 2) one can identify abstract same and abstract different representations in the RMTS task, which are constant even as the perceptual qualities of the underlying object pairs vary.
Figure 5: Linear probing and intervention results. We probe for the intermediate same-different judgments required to perform the RMTS task (blue). We see that probe performance reaches ceiling at layer 5 and maintains throughout the rest of the network. We use the directions defined by the linear probe to intervene on model representations in order to flip an intermediate judgment (green). We see that this intervention succeeds reliably at layer 5 and then drops. We add a vector that is consistent with a pair’s exhibited same-different relation as a control, and see that this has little effect (yellow). The background is colored according to the heatmap in Figure 2(c).
Methods – Patching Novel Representations: In Section 4, we identify independent linear subspaces encoding shape and color information. We now ask: does the content contained in these subspaces matter to the overall same-different computation? One can imagine an ideal version of the same-different relation that is abstracted away from the particular properties of the objects it is comparing. In this setting, the same-different relation could take in object representations where colors and shapes are represented by any arbitrary vector, and the model still correctly judges sameness or difference. To study this, we intervene on either the linear subspaces for shape or color for both objects in a pair, and replace the content found therein with novel representations (See Figure 4(a)). To create a “different” example, we start with a “same” image and replace the shape (or color) representations of both objects with two different novel representations, and to create a “same” example, we start with a “different” image and replace them with two identical novel representations. We then assess whether the model’s decision changes accordingly. We generate novel representations using four methods: 1) we add the representations found within the linear subspaces corresponding to two randomly sampled objects in an IID Validation set, 2) we interpolate between these representations, 3) we sample each dimension randomly from a distribution of embeddings, and 4) we sample each dimension from an OOD random distribution (a mean 0 normal distribution). See Appendix H for technical details.
Results: The results of patching in novel representations are presented in Figure 4(b). Overall, we find the greatest success when patching in added representations, followed by interpolated representations. We observe limited success when patching sampled representations, and no success patching random mean 0 vectors. All interventions perform best in layers 2 and 3, towards the end of the perceptual stage. Overall, this points to a limited form of abstraction present in the relational stage — the same-different operation is not fully abstract (in that it cannot operate over arbitrary vectors), but it does not rely on rote memorization of specific object representations (as it can generalize to additions and interpolations of these representations).
Methods – Linear Interventions: The RMTS task allows us to further characterize the relational stage, as it requires forming an intermediate representation of same and different, and then comparing them. In this analysis, we ask: are these intermediate representations invariant to the perceptual qualities of the object pairs that underlie them? We linearly probe for intermediate same or different judgments from the collection of tokens corresponding to object pairs. The probe consists of a linear transformation mapping the residual stream to 2 dimensions, representing same and different. Each row of this transformation can be viewed as a direction $d$ in the residual stream corresponding to the value that you are probing for (i.e. $d_{same}$ is the linear direction representing $same$). We train one probe for each layer on images from the model’s train set, and test on images from its test set. To understand whether the directions discovered by the probe are causally implicated in model behavior, we (once again) create a counterfactual intervention Nanda et al. (2023). In order to change an intermediate judgment from same to different, we merely add the direction $d_{diff}$ to the intermediate representations of objects that exhibit the same relation, and observe whether the model behaves as if this pair now exhibits the different relation [^6] Prior work normalizes the intervention directions and searches over a scaling parameter. In our setting, we find that simply adding the direction defined by probe weights without rescaling works well, though see Appendix I for further exploration. We use transformerlens for this intervention Nanda & Bloom (2022).. We run this intervention on images from the model’s test set. We also run a control intervention, where we add the incorrect direction (i.e., we add the $d_{same}$ when the object pair already expressed the “same” relation). We anticipate that this will not reliably flip the model’s downstream decisions.
Results: Results of probing and linear interventions are shown in Figure 5. We observe that linear probe performance peaks in the middle layers of the model (layer 5) and then stays high. However, our linear intervention counterfactual accuracy peaks at layer 5 and then drops precipitously. Notably, layer 5 also corresponds to the peak of within-pair attention (See Figure 2. This indicates that, at least in layer 5, there exists a single direction representing same and a single direction representing different. One can flip the intermediate same-different judgment by adding a vector in one of these directions to the residual streams of any pair of objects. Finally, the control intervention completely fails throughout all layers, as expected. Thus, the CLIP-pretrained ViT does in fact generate abstract representations of same and different, and operate on them in order to produce an RMTS response.
Figure 6: We average the best counterfactual intervention accuracy for shape and color and plot it against IID, OOD, and Compositional Test set performance for CLIP, DINO, ImageNet, and from-scratch B/16 models. We observe that increased disentanglement (i.e. higher counterfactual accuracy) correlates with downstream performance. The from-scratch model achieved only chance IID performance in RMTS, so we omitted it from the analysis.
6 Disentanglement Correlates with Generalization Performance
Having object representations that disentangle perceptual properties may enable a model to generalize to out-of-distribution stimuli. As prior work has argued, having disentangled visual representations should enable compositional generalization to unseen combinations of perceptual properties (Higgins et al. (2018); Bengio et al. (2013), cf. Locatello et al. (2019)). To investigate the relationship between disentanglement and generalization, we finetune CLIP, ImageNet, DINO, and randomly-initialized models on a new dataset, where each shape is only ever paired with two distinct colors. We then repeat our analyses in Section 4 to identify independent linear subspaces for shape and color [^7] In order to train these interventions, we must generate counterfactual images. These counterfactual images contain shape-color pairs that are not seen during model training. However, we emphasize that we only use these interventions to derive a disentanglement metric (which we expect to be correlated with compositional generalization), not to train model weights.. We evaluate models in 3 settings: on an IID test set consisting of observed shape-color combinations, on a compositional generalization test set consisting of unobserved shape-color combinations (where each shape and each color have been individually observed), and finally on an OOD consisting of completely novel shapes paired with completely novel colors. We plot the relationship between disentanglement (i.e. counterfactual intervention accuracy) and overall performance on these model evaluations. We find a consistent trend: more disentangled representations correlated with downstream model performance in all cases (See Figure 6).
7 Failure Modes
Previous sections have argued that ViTs trained on same-different tasks implement a two-stage processing pipeline to arrive at a response. In this section, we argue that both perceptual and relational stages could serve as failure points for models, impeding their ability to solve a same-different task. In practice, tasks that rely on relations between objects likely have perceptual and relational stages that are orders of magnitude more complex than those we study here. The results presented herein indicate that solutions targeting either the perceptual (Zeng et al., 2022) or relational (Bugliarello et al., 2023) stages may be insufficient for producing the robust, abstract computations that we desire.
Perceptual Failure Modes: First, we note that models trained from scratch on a discrimination task do not clearly distinguish between perceptual and relational stages (See Figure 2(b)). As Table 1 shows, a model trained on a limited number of shape-color combinations model achieves high train accuracy but poor test accuracy. We hypothesize that a failure to produce object-level representations is precluding the model from solving the discrimination task. To test, we introduce an auxiliary loss to encourage models to form disentangled object-level representations. During training, we introduce token-level probes that are optimized to predict shape information from one linear subspace (i.e., the first 384 dimensions) of the representations generated at an intermediate layer of the model and color information from the complementary linear subspace at that same layer. These probes are optimized during training, and the probing loss is backpropagated through the model. From Table 1, we see that this auxiliary loss enables high test accuracy and also improve compositional generalization.
Aux. Loss Task Patch Size Train Acc. Test Acc. Comp. Acc.
Discrimination 16 77.3% 76.5% 75.9%
Discrimination 16 97.3% 94.6% 86.1%
Discrimination 32 99.1% 88.0% 78.3%
Discrimination 32 99.7% 99.7% 93.5%
Table 1: Adding an auxiliary loss term to encourage disentangled object representations boosts test accuracy and compositional generalization in models trained from scratch on the discrimination task.
Relational Failure Modes: We might ask whether the same approach can help models trained from-scratch on the RMTS task – models that achieve only chance performance. Despite our auxiliary loss probes achieving high IID probing accuracy, our intervention still results in chance test accuracy (See Appendix J). This provides evidence that 1) models may fail at either the perceptual or relational stages and 2) that the relation defined by the discrimination task is simple to solve with disentangled object representations, while the relation defined in the RMTS task is still far from trivial.
8 Discussion
Related Work: This work takes inspiration from the field of mechanistic interpretability, which seeks to characterize the algorithms that neural networks implement (Olah, 2022). Though many of these ideas originated in the domain of NLP (Wang et al., 2022; Hanna et al., 2024; Feng & Steinhardt, 2023; Wu et al., 2024c; Merullo et al., 2023; Geva et al., 2022; Meng et al., 2022) and in toy settings (Nanda et al., 2022; Elhage et al., 2022; Li et al., 2022), they are beginning to find applications in computer vision (Fel et al., 2023; Vilas et al., 2024; Palit et al., 2023). These techniques augment an already-robust suite of tools that visualize the features (rather than algorithms) that vision models use (Olah et al., 2017; Selvaraju et al., 2017; Simonyan et al., 2014). Finally, this study contributes to a growing literature employing mechanistic interpretability to address debates within cognitive science (Millière & Buckner, 2024; Lepori et al., 2023a; Kallini et al., 2024; Traylor et al., 2024).
Conclusion: The ability to compute abstract visual relations is a fundamental aspect of biological visual intelligence and a crucial stepping stone toward useful and robust artificial vision systems. In this work, we demonstrate that some finetuned ViTs adopt a two-stage processing pipeline to solve same-different tasks – despite having no obvious inductive biases towards this algorithm. First, models produce disentangled object representations in a perceptual stage; then models compute a somewhat abstract version of the same-different computation in a relational stage. Finally, we observed a correlation between disentanglement and generalization and noted that models might fail to learn either the perceptual or relational operations necessary to solve a task. One question raised by this study is why CLIP-pretrained models perform favorably (and adopt this two-algorithm so cleanly) relative to other pretrained models. Raghu et al. (2021) finds that models pretrained on more data tend to learn local attention patterns in early layers, followed by global patterns in later layers. This might give CLIP a particular incentive to create local object representations, which are then used in relational operations. Future work might test this hypothesis. Future work might also focus on pinning down the precise mechanisms (i.e., the attention heads and MLPs) used to implement the perceptual and relational stages or on generalizing these findings to more complex datasets.
Appendix A Constructing the Objects
Figure A demonstrates a single instance of the $16$ shapes and $16$ colors used in our datasets. Any shape can be paired with any color to create one of $256$ unique objects. Note that object colors are not uniform within a given object. Instead, each color is defined by three different Gaussian distributions—one for each RGB channel in the image—that determine the value of each object pixel. For example, the color red is created by these three distributions: $\mathcal{N}(\mu=233,\sigma=10)$ in the red channel, $\mathcal{N}(\mu=30,\sigma=10)$ in the green channel, and $\mathcal{N}(\mu=90,\sigma=10)$ in the blue channel. All color distributions have a variance fixed at $10$ to give them an equal degree of noise. Any sampled values that lie outside of the valid RGB range of $[0,255]$ are clipped to either 00 or $255$. Object colors are re-randomized for every image, so no two objects have the same pixel values even if they are the same color. This was done to prevent the models from learning simple heuristics like comparing single pixels in each object.
Figure 7: All 16 unique shapes and colors used to construct the Discrimination and RMTS tasks.
Appendix B ViT B/16 All Model Results
See Table 2 for behavioral results from all models trained on discrimination and RMTS tasks.
Pretrain Task Train Acc. IID Test Acc.
CLIP Disc. 100% 99.3%
CLIP RMTS 100% 98.3%
ImageNet Disc. 100% 97.5%
ImageNet RMTS 99.7%89.3%
DINO Disc. 100% 95.6%
DINO RMTS 100% 87.7%
Disc. 95.9%80.5%
RMTS 54.7% 49.6%
Table 2: All behavioral results for ViT-B/16 models.
Appendix C CLIP ViT-b32 Model Analyses
C.1 Attention Pattern Analysis
See Figure 8.
Figure 8: CLIP B/32 attention pattern analysis.
C.2 Perceptual Stage Analysis
See Figure 9.
Figure 9: CLIP B/32 perceptual stage analysis.
C.3 Relational Stage Analysis
See Figure 10.
Figure 10: CLIP B/32 relational stage analysis.
C.4 Generalization Results
We present CLIP-B/32 model results for models finetuned on all shape-color combinations, as well as only 32 shape-color combinations, as in Section 6. We present compositional generalization accuracy (when applicable) as well as OOD generalization accuracy. We find that all models perform quite well in-distribution and under compositional generalization. Accuracy drops somewhat for RMTS OOD stimuli. All results are in presented in Table 3.
Task # Combinations Seen Train Acc. IID Test Acc. Comp. Gen. Acc. OOD Acc.
Disc. All 100% 99.6% N/A 95.8%
Disc. 32 99.7% 98.5% 98.5% 98.0%
RMTS All 100% 97.4% N/A 90.5%
RMTS 32 100% 98.3% 98.3% 86.2%
Table 3: All behavioral results for CLIP-B/32 models.
Appendix D Attention Patterns Examples
See Figure 11.
Figure 11: Example attention head patterns for CLIP ViT-B/16 and From Scratch ViT-B/16 trained on the Discrimination task.
Appendix E Distributed Alignment Search Technical Details
Approach
We apply a form of distributed alignment search (Geiger et al., 2024) in order to assess whether the object representations formed by the perceptual stage of ViTs are disentangled with respect to shape and color (the two axes of variation present in our dataset). For ViT B/32 models, each object is contained within the bounds of a single patch, making DAS straightforward to run: we train an intervention over model representations corresponding to the patches containing individual objects that we wish to use as source and base tokens. For ViT B/16 models, each object is contained in four patches. Here, we train a single intervention that is shared between all four patches comprising the base and source objects. Importantly, because we wish to isolate whole object representations, rather than patch representations, we randomly shuffle the 4 patches comprising the source object before patching information into the base object. For example, the top-right patch of the base object might be injected with information from the bottom-left patch of the source object. This intervention should only succeed if the model contains a disentangled representation of the whole object, and if this representation is present in all four patches comprising that object. Given these stringent conditions, it is all the more surprising that DAS succeeds. During test, we intervene in a patch-aligned manner: The vector patched into the top-right corner of the base image representation is extracted from the top-right corner of the source image.
Data
To train the DAS intervention, we must generate counterfactual datasets for every subspace that we wish to isolate. To generate a discrimination dataset that will be used to identify a color subspace, for example, we find examples in the model’s training set where objects only differ along the color dimension (e.g., object1 expresses color1, object2 expressed color2. We randomly select one object to intervene on and generate a counterfactual image. WLOG consider intervening on object1. Our counterfactual image contains one object (the counterfactual object) that expresses color2. Our intervention is optimized to extract color information from the counterfactual object and inject it into object1, changing the model’s overall discrimination judgment from different to same. Importantly, the counterfactual image label is also different. Thus, our intervention is designed to work only if the intervention transfers color information. We follow a similar procedure for to generate counterfactuals that can be used to turn a same image into a different image. In this case, both base and counterfactual images are labelled same, but the counterfactual same image contains objects that are a different color than those in the base image. The counterfactual color is patched into one of the objects in the base image, rendering the objects in the base image different along the color axis.
For the RMTS DAS dataset, we generate counterfactuals similarly to the discrimination dataset. We select a pair of objects randomly (except either the display pair or sample pair). We then choose the source object in the other pair. We edit the color or shape property of just this source object, and use this as the counterfactual. For these datasets, the data is balanced such that 50% of overall labels are changed from same to different, but also 50% of intermediate pair labels are changed from same to different. Note that flipping one intermediate label necessarily flips the hierarchical label. Thus, if the source object is in a pair expressing the same relationship, then the counterfactual image will have the opposite label as the base image before intervention. In these cases, the intervention could succeed by transferring the hierarchical image label, rather than by transferring particular color or shape properties from one object to another. However, this only occurs approximately 50% of the time. That is because it occurs in 100% of samples when both pairs exhibit same, which occurs 25% of the time (half of the hierarchical same images), and 50% of the time when one pair exhibits same and the other exhibits different, which occurs 50% of the time (all of the hierarchical different images). However, this strategy provides exactly the incorrect incorrect result in the other 50% of cases. Nonetheless, this behavior might explain why RMTS DAS results maintain at around 50% deeper into the model.
For all datasets, we generate train counterfactual pairs from the model train split, validation pairs from the validation split, and test pairs from the model test split. We generate 2,000 counterfactual pairs for both splits. Note that in the case of models trained in the compositional generalization experiments (i.e. those found in Section 6), the counterfactual image may contain shape-color pairs that were not observed during training. However, training our interventions has no bearing on the model’s downstream performance on held-out data, though correlation between disentanglement and compositional generalization is thus not extremely surprising. See Figure 12 for examples of counterfactual pairs used to train interventions.
Refer to captionRefer to caption
(a) Discrimination: Color counterfactual pair. The brown color from the object on the right will be patched into the orange color in the object on the left.
Refer to captionRefer to caption
(b) RMTS: Color counterfactual pair. The dark blue color from the display pair object on the right will be patched into one of the green sample objects on the left.
Refer to captionRefer to caption
(c) Discrimination: Shape counterfactual pair. The two-circles shape from the object on the right will be patched into the cross shape in the object on the left.
Refer to captionRefer to caption
(d) RMTS: Shape counterfactual pair. The thin star shape from the sample pair object on the right will be patched into one of the orange objects on the left.
Figure 12: Counterfactual pairs used to train DAS interventions.
Intervention Details
DAS requires optimizing 1) a rotation matrix over representations and 2) some means of identifying appropriate dimensions over which to intervene (Geiger et al., 2024). Prior work has largely heuristically selected contiguous subspaces over which to intervene (Geiger et al., 2024; Wu et al., 2024c). In this work, we relax this heuristic, identifying dimensions by optimizing a binary mask over model representations as we optimize the rotation matrix (Wu et al., 2024a). We follow best practices from differentiable pruning methods like continuous sparsification (Savarese et al., 2020), annealing a sigmoid mask into a binary mask over the course of training, using an exponential temperature scheduler. We also introduce an L0 penalty to encourage sparse masking. We use default parameters suggested by the pyvene library for Boundless DAS, another DAS method that optimizes the dimensions over which to intervne. Our rotation matrix learning rate is 0.001, our mask learning rate is 0.01, and we train for 20 epochs for each subspace, independently for each model layer. We add a scalar multiplier of 0.001 to our L0 loss term, which balances the magnitude of L0 loss with the normal cross entropy loss that we are computing to optimize the intervention. Our temperature is annealed down to 0.005 over the course of training, and then snapped to binary during testing. Finally, we optimize our interventions using the Adam optimizer (Kingma & Ba, 2014). These parameters reflect standard practice for differentiable masking for interpretability (Lepori et al., 2023b).
Appendix F Perceptual Stage Analysis Controls
As a control for our DAS analysis presented in Section 4, we attempt to intervene using the incorrect source token in the counterfactual image. If this intervention fails, then it provides evidence that the information transferred in the standard DAS experiment is actually indicative of disentangled local object representations, rather than information that may be distributed across all objects. We note that this control could succeed at flipping same judgments to different, but will completely fail in the opposite direction. As shown in Figure 13, these controls do reliably fail to achieve above-chance counterfactual intervention accuracy.
Figure 13: Controls for DAS Analysis (Section 4). (a) We intervene on objects using the irrelevant source token in the counterfactual stimuli. For example, if the DAS interventions are meaningful, patching a bowtie shape into the base image in the top row should not change the model’s decision to “same.” (b) The controls fail to flip CLIP ViT’s decisions at a rate above chance accuracy, indicating that the DAS results presented in Section 4 are indeed the result of meaningful interventions.
Appendix G Perceptual Stage Analysis: Other Models
See Figures 14,15,16 for DINO, ImageNet, and From Scratch DAS results.
Figure 14: DAS results for DINO ViT-B/16.
Figure 15: DAS results for ImageNet ViT-B/16.
Figure 16: DAS results for From Scratch ViT-B/16.
Appendix H Novel Representations Analysis Technical Details
During the novel representations analysis of Section 5, we patch in novel vectors into the subspaces identified using DAS. Notably, we must patch into the {color, shape} subspace for both objects that we wish to intervene on, rather than just one. This is because we want to analyze whether the same-different relation can generalize to novel representations. For example, if two objects in a discrimination example share the same shape, but one is blue and one is red , we would like to know whether we can intervene to make the color property of each object an identical, novel vector, such that the model’s decision will flip from different to same. We run this analysis on the IID test set of the DAS data.
We create these vectors using four different methods. For these methods, we first save the embeddings found in the subspaces identified by DAS for all images in the DAS Validation set.
  1. 1.
    Addition: We sample two objects in the validation set, and add their subspace embeddings. For ViT-B/16, we patch the resulting vector in a patch-aligned manner: The vector patched into the top-right corner of the base image representation is generated by adding the top-right corners of each embedding found within the subspace of the sampled validation images.
  2. 2.
    Interpolation: Same as Method 1, except vectors are averaged dimension-wise.
  3. 3.
    Sampled: We form one Gaussian distribution per embedding dimension using our saved validation set embeddings. We independently sample from these distributions to generate a vector that is patched into the base image. This single vector is patched into all four object patches for ViT-B/16.
  4. 4.
    Random Gaussian: We randomly sample from a normal distribution with mean 0 and standard deviation 1 and patch that into the base image. This single vector is patched into all four object patches for ViT-B/16.
Appendix I Relational Stage Analysis: Other Models
I.1 Novel Representation Analysis
See Figures 18,19,20,21,22 for DINO Discrimination/RMTS, ImageNet Discrimination/RMTS, and From Scratch Discrimination Novel Representation Analysis results.
Figure 17: Novel Representation Analysis for CLIP ViT-B/16 (RMTS).
Figure 18: Novel Representation Analysis for DINO ViT-B/16 (Disc.).
Figure 19: Novel Representation Analysis for DINO ViT-B/16 (RMTS).
Figure 20: Novel Representation Analysis for ImageNet ViT-B/16 (Disc.).
Figure 21: Novel Representation Analysis for ImageNet ViT-B/16 (RMTS).
Figure 22: Novel Representation Analysis for From Scratch ViT-B/16 (Disc.).
I.2 Abstract Representations of Same and Different
We run the linear probe and linear intervention analysis from Section 5 on DINO B/16 and ImageNet B/16. We find that the intervention works much less well on these models than on CLIP B/16 or B/32. This indicates that these models are not using one abstract representation of same and one representation of different that is agnostic to perceptual qualities of the input image.
Additionally, we try to scale the directions that we are adding to the intermediate representations by 0.5 and 2.0, and find that neither of these versions of the intervention work much better well for either model. See Figure 23 and 24 for DINO and ImageNet results.
Figure 23: Scaled linear probe & intervention analysis for DINO ViT-B/16.
Figure 24: Scaled linear probe & intervention analysis for ImageNet ViT-B/16.
Appendix J Relational Failure Mode
Adding an auxiliary loss term to encourage disentangled object representations does not help for RMTS (See 4).
Aux. Loss Task Patch Size Train Acc. Test Acc. Comp. Acc.
RMTS 16 49.2% 50.1% 50.0%
RMTS 16 53.9% 54.4% 54.1%
RMTS 32 87.7% 51.4% 51.4%
RMTS 32 70.2% 50.2% 50.2%
Table 4: Adding an auxiliary loss term to encourage disentangled object representations does not help for RMTS.
Appendix K Compute Resources
We employed compute resources at a large academic institution. We scheduled jobs with SLURM. Finetuning models on these relational reasoning tasks using geforce3090 GPUs required approximately 200 GPU-hours of model training. Running DAS over each layer in a model required approximately 250 GPU-hours. The remaining analyses took considerably less time, approximately 50 GPU-hours in total. Preliminary analysis and hyperparameter tuning took considerably more time, approximately 2,000 GPU-hours in total. The full research project required approximately 2,500 GPU-hours.
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: See Section 1.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: See Section 8.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification: [N/A]
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: Sections 2,3,4, 5, Appendices EH
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: We link to an anonymized repository containing our code.
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  25. 6.
    Experimental Setting/Details
  26. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  27. Answer: [Yes]
  28. Justification: Sections 2,3,4, 5, Appendices EH
  29. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in appendix, or as supplemental material.
  30. 7.
    Experiment Statistical Significance
  31. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  32. Answer: [No]
  33. Justification: Our results primarily concern analyses conducted on finetuned versions of pretrained models. We present model analyses separately so as not to lose information through aggregation, at the expense of reporting statistical significance.
  34. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  35. 8.
    Experiments Compute Resources
  36. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  37. Answer: [Yes]
  38. Justification: See Appendix K.
  39. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  40. 9.
    Code Of Ethics
  41. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  42. Answer: [Yes]
  43. Justification: [N/A]
  44. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  45. 10.
    Broader Impacts
  46. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  47. Answer: [N/A]
  48. Justification: This work centers around issues at the intersection of cognitive science and deep learning. This work does not contribute toward practical algorithms or techniques that will see broad deployment.
  49. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  50. 11.
    Safeguards
  51. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  52. Answer: [N/A]
  53. Justification: The models used in this paper are standard vision transformers finetuned on a simple synthetic task. There are no such risks.
  54. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  55. 12.
    Licenses for existing assets
  56. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  57. Answer: [Yes]
  58. Justification: [N/A]
  59. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  60. 13.
    New Assets
  61. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  62. Answer: [N/A]
  63. Justification: [N/A]
  64. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  65. 14.
    Crowdsourcing and Research with Human Subjects
  66. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  67. Answer: [N/A]
  68. Justification: [N/A]
  69. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  70. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  71. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  72. Answer: [N/A]
  73. Justification: [N/A]
  74. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.